# A Comprehensive Evaluation on Event Reasoning of Large Language Models

### v1

@@ 5.2 To what extent do LLMs have the event schema knowledge?
We evaluate CEC and CRR on the schema level. The results are shown in Figure 2, and the average scores are reported in Table 4. We find LLMs already have event schema knowledge and can complete both CEC and CRR tasks at the schema level to some extent. However, in Table 4, we observe that S-CEC lags I-CEC, suggesting that LLMs are more adept at reasoning at the instance level.
@@
We probe how event schema knowledge increases with the development of LLMs. We depict CEC performance comparisons of LLMs on instance- and schema-level in Figure 4. When the models initially can reason about events, they also have event schema knowledge. At this time, models can perform comparatively or even better in schema-level event reasoning. With the development, models perform instance-level reasoning better than schema-level. It indicates that the accumulation of event schema knowledge falls behind the reasoning at the instance level. This finding demonstrates that enhancing event schema knowledge may further improve these abilities to obtain better general LLMs.
@@ 5.3 Are LLMs aligned with humans in the aspect of leveraging event schema knowledge
However, we find the LLMs would generate event types of different words but correct contents. To mitigate this problem, we prepare a list of candidate event types for each data to make it a classification setting. To keep the task difficult, we first conduct KMeans clustering on all event types in our dataset444We use all-mpnet-base-v2 for encoding.. We obtain 1000 clusters. For each data, we assign 20 random candidates in total including the correct ones. The negative event types are chosen from different clusters.
After the generation, we calculate the accuracy of event types and F1-scores of relation triplets respectively comparing with the human-labeled event schema. We regard a correct triplet if all the head and tail event types and the inter-relation align with the human labels. We show detailed examples in Figures 9-10 in Appendix.
The results are in Table 5. We find only GPT4 can generate correct event types while other models all fail. For relation triplet generation, even GPT4 can not output proper event schemas555GPT4 excels other may be attributed to 1) its better alignment. 2) The dataset is originally generated by GPT4.. It significantly suggests that LLMs may not leverage event schema knowledge as humans when solving event reasoning tasks. Alignment of using such knowledge could further improve the performances.
| CEC | CRR | |||
| ET | Rel | ET | Rel | |
| GPT4 | 70.71 | 37.30 | 70.66 | 49.41 |
| GPT3.5 | 13.43 | 15.78 | 18.55 | 21.14 |
| Mistral-7B | 11.15 | 9.00 | 11.88 | 15.15 |
@@ 5.4 Can LLMs perform better event reasoning with explicit guidance of leveraging event schema knowledge?
We design two guiding methods:
Direct: Directly add the event type of each instance event into the prompt.
CoT: Guide the LLMs in a CoT-style to 1) generate the event types of each instance event. 2) reason with the event types. This is a more practical method since we would not know the event types in advance in real scenarios.
| CEC | CRR | |||
| w.o.S | w.t.S | w.o.S | w.t.S | |
| GPT4 | 63.80 | 69.89 (6.09) | 61.2 | 63.11 (1.91) |
| GPT3.5 | 50.18 | 60.92 (10.74) | 39.52 | 45.99 (6.47) |
| Mistral-7B | 57.71 | 63.26 (5.55) | 44.00 | 47.07 (3.07) |
| Llama2-7B | 30.29 | 38.17 (7.88) | 34.00 | 43.35 (9.35) |
| WizardLM | 33.69 | 29.93 (3.76) | 37.00 | 44.91 (7.91) |
| Vicuna-7B | 31.18 | 34.41 (3.23) | 42.00 | 42.40 (0.40) |
We show the performances of direct guidance in Table 6. We find incorporating event schema knowledge significantly improves event reasoning. It shows great potential to solve event reasoning with the fusion of event schema knowledge.
We report the results of the CoT guidance in Table 7. We only report results of GPT-4 since we find other models are unable to follow this instruction. We find in CEC, CoT guidance can improve performance. However, the improvement of CoT lags those of Direct, indicating great space for better methods. Developing advanced guidance for all LLMs remains a challenging research problem. We show the example of prompt and GPT4 generation in Figures 11 and 12 in Appendix.
| CEC | w.o.S | w.t.S |
| GPT4 | 63.80 | 67.92 (4.12) |
@@ 6 Related Work
Du et al. (2022) aims to select the accurate cause or effect event from candidates. Zhou et al. (2019) serves as a dataset for event temporal reasoning. Current works present a scenario of incorporating counterfactual reasoning Qin et al. (2019, 2020). In addition to single-event relation reasoning, existing works also reason events according to diversified event relations Poria et al. (2021); Han et al. (2021); Yang et al. (2022). Tao et al. (2023b) further unifies datasets of several event-inter relations to transfer event relational knowledge to unseen tasks.
@@ 7 Conclusion
In this paper, we evaluate the event reasoning of LLMs. We introduce a novel benchmark which features both levels of schema and instance. It evaluates event schema knowledge and reasoning abilities. Besides, can be used to comprehensively evaluate the event reasoning in various relations and reasoning paradigms. We conduct extensive experiments on . We obtain many insights such as: 1) LLMs have the abilities of event reasoning, but are far from satisfactory and are unbalanced in different relations and reasoning paradigms. 2) LLMs have event schema knowledge. However, with the development of LLMs, this knowledge increases slowly compared with the increase of abilities of event instance reasoning. 3) LLMs are not aligned with human to leaverage event schema knowledge in event reasoning. 4) Based on the findings, we design two methods, namely Direct and CoT, to guide the LLMs to utilize event schema knowledge. With our designed guidances for utilizing event schema knowledge, LLMs can perform better event reasoning.
Limitations
We guide the LLMs to utilize the event schema knowledge in two ways. The Direct effects most. However, the more practical way CoT falls behind Direct indicating the potential of a better method of guidance. We leave it to future work.

### v2

@@ 5.2 To what extent do LLMs have the event schema knowledge?
We evaluate CEC and CRR on the schema level. The results are shown in Figure 2 and detailed results in the Appendix, and the average scores are reported in Table 4. We find LLMs already have event schema knowledge and can complete both CEC and CRR tasks at the schema level to some extent. However, in Table 4, we observe that S-CEC lags I-CEC, suggesting that LLMs are more adept at reasoning at the instance level.
@@
We probe how event schema knowledge increases with the development of LLMs. We depict CEC and CRR performance comparisons of LLMs on instance- and schema-level in Figure 4. When the models initially can reason about events, they also have event schema knowledge. At this time, models can perform comparatively or even better in schema-level event reasoning. With the development, models perform instance-level reasoning on par with schema-level. It indicates that the accumulation of event schema knowledge falls behind the reasoning at the instance level. This finding demonstrates that enhancing event schema knowledge may further improve these abilities to obtain better general LLMs.
@@ 5.3 Are LLMs aligned with humans in the aspect of leveraging event schema knowledge
However, we find the LLMs would generate event types of different words but correct contents. To mitigate this problem, we prepare a list of candidate event types for each data to make it a classification setting. To keep the task difficult, we first conduct KMeans clustering on all event types in our dataset555We use all-mpnet-base-v2 for encoding.. We obtain 1000 clusters. For each data, we assign 20 random candidates in total including the correct ones. The negative event types are chosen from different clusters.
After the generation, we calculate the accuracy of event types and F1-scores of relation triplets respectively comparing with the human-labeled event schema. We regard a correct triplet if all the head and tail event types and the inter-relation align with the human labels. We show detailed examples in the Appendix.
The results are in Table 5. We find only GPT4 and GPT4o can generate correct event types. GPT4 excels may be attributed to 1) its better alignment. 2) The dataset is originally generated by GPT4. Therefore in this part, we mainly evaluate other models rather than GPT4 series. However, we find rest models all fail to generate corresponding schema knowledge. For relation triplet generation, even GPT4 can not output proper event schemas. It significantly suggests that LLMs may not leverage event schema knowledge as humans when solving event reasoning tasks. Alignment of using such knowledge could further improve the performances.
@@ 5.4 Can LLMs perform better event reasoning with explicit guidance of leveraging event schema knowledge?
We conduct experiments in which we directly add the schema event of each instance event in the question into the prompt. We also add relations of these schema events.
We demonstrate the performances of this guidance in Table 6 and 7. We find incorporating event schema knowledge significantly improves event reasoning. It shows great potential to solve event reasoning with the fusion of event schema knowledge. These results provide important insights that event schema knowledge could be used as memory to improve solving practical and domain-specific problems. In this view, constructing and retrieving proper event schema knowledge is another challenge. We leave them to future works.
@@ 6 Related Work
Du et al. (2022a) aims to select the accurate cause or effect event from candidates. Zhou et al. (2019) serves as a dataset for event temporal reasoning. Current works present a scenario of incorporating counterfactual reasoning Qin et al. (2019, 2020). In addition to single-event relation reasoning, existing works also reason events according to diversified event relations Poria et al. (2021); Han et al. (2021); Yang et al. (2022). Tao et al. (2023b) further unifies datasets of several event-inter relations to transfer event relational knowledge to unseen tasks.
@@ 7 Conclusion
In this paper, we evaluate the event reasoning of LLMs. We introduce a novel benchmark which features both levels of schema and instance. It evaluates event schema knowledge and reasoning abilities. Besides, can be used to comprehensively evaluate the event reasoning in various relations and reasoning paradigms. We conduct extensive experiments on . We obtain many insights such as: 1) LLMs have the abilities of event reasoning, but are far from satisfactory and are unbalanced in different relations and reasoning paradigms. 2) LLMs have a comprehension of event schema knowledge. 3) LLMs are not aligned with human to leaverage event schema knowledge in event reasoning. 4) Based on the findings, we guide the LLMs to utilize event schema knowledge. With our guidance, LLMs can perform better event reasoning.
@@
Detailed Results
We show detailed results in Table 8 (S-CEC), Table 9 (I-CEC), Table 10 (S-CRR), and Table 10 (I-CRR).
Model | Before | After | Cause | Result | isSubevent | hasSubevent | Temporal | Causal | Hierarchy | All |
GPT4o | 66.67 | 53.57 | 68.52 | 71.98 | 76.67 | 82.35 | 57.72 | 71.33 | 77.92 | 68.93 |
GPT4 | 64.1 | 53.57 | 66.67 | 72.84 | 71.67 | 76.47 | 56.91 | 71.68 | 72.73 | 68.11 |
GPT3.5 | 46.15 | 48.81 | 74.07 | 72.41 | 65.00 | 70.59 | 47.97 | 72.73 | 66.23 | 65.43 |
Mistral-7B | 38.46 | 47.62 | 44.44 | 57.33 | 56.67 | 52.94 | 44.72 | 54.90 | 55.84 | 52.47 |
Qwen2-7B | 56.41 | 50.00 | 57.41 | 68.53 | 63.33 | 58.82 | 52.03 | 66.43 | 62.34 | 62.14 |
Baichuan2-7B | 51.28 | 48.81 | 53.7 | 55.17 | 48.33 | 58.82 | 49.59 | 54.90 | 50.65 | 52.88 |
Orca2-7B | 53.85 | 42.86 | 64.81 | 67.67 | 55.00 | 52.94 | 46.34 | 67.13 | 54.55 | 59.88 |
Chatglm2-6B | 38.46 | 45.24 | 57.41 | 61.21 | 55.00 | 70.59 | 43.09 | 60.49 | 58.44 | 55.76 |
Interlm2-7B | 53.85 | 52.38 | 72.22 | 69.83 | 71.67 | 64.71 | 52.85 | 70.28 | 70.13 | 65.84 |
Llama2-7B | 33.33 | 36.90 | 50.00 | 48.71 | 41.67 | 58.82 | 35.77 | 48.95 | 45.45 | 45.06 |
Vicuna-7b | 15.38 | 28.57 | 16.67 | 28.02 | 33.33 | 11.76 | 24.39 | 25.87 | 28.57 | 25.93 |
Model | Before | After | Cause | Result | isSubevent | hasSubevent | Temporal | Causal | Hierarchy | All |
GPT4o | 60.00 | 52.38 | 69.09 | 70.64 | 75.00 | 58.82 | 54.84 | 70.34 | 71.43 | 66.60 |
GPT4 | 70.00 | 58.33 | 69.09 | 69.36 | 76.67 | 70.59 | 62.1 | 69.31 | 75.32 | 68.43 |
GPT3.5 | 50.00 | 44.05 | 74.55 | 71.06 | 73.33 | 52.94 | 45.97 | 71.72 | 68.83 | 64.77 |
Mistral-7B | 45.00 | 35.71 | 63.64 | 62.98 | 45.00 | 47.06 | 38.71 | 63.10 | 45.45 | 54.18 |
Qwen2-7B | 55.00 | 52.38 | 76.36 | 72.34 | 50.00 | 29.41 | 53.23 | 73.10 | 45.45 | 63.75 |
Baichuan2-7B | 22.5 | 26.19 | 43.64 | 29.36 | 30.00 | 29.41 | 25.00 | 32.07 | 29.87 | 29.94 |
Orca2-7B | 55.00 | 47.62 | 56.36 | 67.66 | 55.00 | 58.82 | 50.00 | 65.52 | 55.84 | 60.08 |
Chatglm2-6B | 32.50 | 9.52 | 50.91 | 33.19 | 33.33 | 29.41 | 16.94 | 36.55 | 32.47 | 30.96 |
Interlm2-7B | 57.50 | 40.48 | 67.27 | 68.09 | 68.33 | 58.82 | 45.97 | 67.93 | 66.23 | 62.12 |
Llama2-7B | 32.5 | 28.57 | 45.45 | 25.53 | 33.33 | 23.53 | 29.84 | 29.31 | 31.17 | 29.74 |
Vicuna-7b | 22.5 | 29.76 | 20.00 | 28.94 | 25.00 | 29.41 | 27.42 | 27.24 | 25.97 | 27.09 |
Model | Temporal | Causal | Hierarchy | All |
GPT4o | 51.89 | 66.98 | 60.00 | 62.05 |
GPT4 | 61.08 | 66.28 | 53.91 | 63.01 |
GPT3.5 | 46.49 | 58.60 | 52.17 | 54.52 |
Mistral-7B | 46.49 | 55.58 | 45.22 | 51.64 |
Qwen2-7B | 41.62 | 58.14 | 46.09 | 52.05 |
Baichuan2-7B | 49.73 | 53.49 | 47.83 | 51.64 |
Orca2-7B | 33.51 | 52.79 | 41.74 | 46.16 |
Chatglm2-6B | 50.27 | 56.51 | 40.87 | 52.47 |
Interlm2-7B | 45.41 | 51.63 | 42.61 | 48.63 |
Llama2-7B | 35.14 | 51.40 | 46.96 | 46.58 |
Vicuna-7B | 47.57 | 53.95 | 52.17 | 52.05 |
Model | Temporal | Causal | Hierarchy | All |
GPT4o | 48.92 | 67.74 | 61.74 | 62.04 |
GPT4 | 52.15 | 67.05 | 70.43 | 63.81 |
GPT3.5 | 36.56 | 46.54 | 46.09 | 43.95 |
Mistral-7B | 43.55 | 62.44 | 46.96 | 55.24 |
Qwen2-7B | 36.56 | 53.69 | 42.61 | 47.62 |
Baichuan2-7B | 36.02 | 51.38 | 37.39 | 45.31 |
Orca2-7B | 41.94 | 49.77 | 33.04 | 45.17 |
Chatglm2-6B | 46.77 | 54.15 | 37.39 | 49.66 |
Interlm2-7B | 43.55 | 61.29 | 64.35 | 57.28 |
Llama2-7B | 35.48 | 43.55 | 41.74 | 41.22 |
Vicuna-7b | 44.62 | 54.15 | 52.17 | 51.43 |
Model | Version | URL |
GPT4o | gpt-4o-2024-05-13 | |
GPT4 Achiam et al. (2023) | gpt-4-0125-preview | |
GPT3.5 | gpt-3.5-turbo-1106 | |
Mistral-7B Jiang et al. (2023) | mistralai/Mistral-7B-Instruct-v0.2 | https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2 |
Qwen2-7B Yang et al. (2024) | Qwen/Qwen2-7B-Instruct | https://huggingface.co/Qwen/Qwen2-7B-Instruct |
Baichuan2-7B Yang et al. (2023) | baichuan-inc/Baichuan2-7B-Chat | https://huggingface.co/baichuan-inc/Baichuan2-7B-Chat |
Orca2-7B Mitra et al. (2023) | microsoft/Orca-2-7b | https://huggingface.co/microsoft/Orca-2-7b |
Chatglm2-7B GLM et al. (2024) | THUDM/chatglm2-6b | https://huggingface.co/THUDM/chatglm2-6b |
Internlm2-7B Cai et al. (2024) | internlm/internlm2_5-7b-chat | https://huggingface.co/internlm/internlm2_5-7b-chat |
Llama2-7B Touvron et al. (2023) | NousResearch/Llama-2-7b-chat-hf | https://huggingface.co/NousResearch/Llama-2-7b-chat-hf |
Vicuna-7B Chiang et al. (2023) | lmsys/vicuna-7b-v1.3 | https://huggingface.co/lmsys/vicuna-7b-v1.3 |
Annotation Process
In this section, we describe the annotation of the third stage in detail. We illustrate the annotation platform in Figure 5. After logining in and selecting assigned data (i.e. the prompting graphs and prepared candidate events), to annotate this data, the annotator should undergo three steps:
1)
Modify the schema graph As shown in Figure 5, Target Events are the queried event, answer event, and their relation. Context Events are other events in schema graph. The Event Schema Graph shows the queried event in red, answer event in blue, and other context event in green. The first step is to rewrite the events and relations in in schema graph to make them strictly valid.
2)
Modify the instance graph The Event Instances are the instance events of the events in the schema graph. The second step is to rewrite the instance events to make the instance graph valid and is relationally consistent with the schema graph.
3)
Modify candidate events The Candidates are the events in both schema and instance levels. They are the candidates of the answer event in the final CEC question. The last step is to modify and select candidates as negative choices considering the answer event where each candidate event consists of a schema and an instance event.
Model Details
We list all model details in Table 12.